The board of examiners that administers the real estate brokerʹs examination in a certain state found that the mean score on the test was 375 and the standard deviation was 36. If the board wants to set the passing score so that only the best 80% of all applicants pass, what is the passing score? Assume that the scores are normally distributed
Answer #1

solution

Using standard normal table,

P(Z > z) = 80%

= 1 - P(Z < z) = 0.80

= P(Z < z ) = 1 - 0.80

= P(Z < z ) = 0.20

= P(Z <-0.84 ) = 0.20  

z = -0.84 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x=-0.84 *36+375

x= 344.76